OpenStudy (anonymous):

when f(x)=3x^2 and g(x)=x-5 what is the value of f(g(3))

OpenStudy (anonymous):

see first find g(3) and then use the result in f(x)

OpenStudy (anonymous):

So g(3)=3-5 g(3)=-2 f(-2)=3(-2)^2 f(-2)=12

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

welcome dear...